# Scalable hybrid quantum Monte Carlo simulation of U(1) gauge field coupled to fermions on GPU

### v3

@@ Introduction
The emergence of Dirac spin liquid (DSL), i.e. an emergent $U(1)$ gauge field coupled to gapless relativistic spinons in (2+1)d [Hermele2004, Assaad2005, Hermele2005, ranProjected2007, ranSpontaneous2009, XYXuU12018, wangDynamics2019, LuoTwistedDSL22, WietekQED3, SeifertSpinPeierls24, janssenConfinement2020], or, in high-energy physics terms, a deconfined phase in 3D quantum electrodynamics (QED3) with $N_{f}$ flavors of massless fermions [Fiebig1990, Herbut2003, Fiore2005, Armour2011, chester2016towards, di2017scaling, karthikNumerical2019, Albayrak2022], has attracted continuous attention from condensed matter and high-energy physicists over the past decades. For sufficiently large $N_{f}$, QED3 is expected to flow to a conformal field theory (CFT) in the infrared [karthikNumerical2019, chester2016towards], containing a variety of scaling operators including fermion bilinear currents, mass operators, and monopole operators. Such a critical state has been argued to emerge in frustrated spin systems, where no magnetic order appears and the spinons acquire Dirac dispersions, and their interactions are mediated by fluctuating $U(1)$ gauge fields [Hastings2003, Hermele2004, Assaad2005, Hermele2005, ranProjected2007, ranSpontaneous2009, Iqbal2011, Imada2014, White2015, Hu2015, Iqbal2016, He2017, Gong2017, Saadamand2017, XYXuU12018, wangDynamics2019, song19, He2019, Ferrari2019, nomuraDirac2021, LuoTwistedDSL22, WietekQED3, SeifertSpinPeierls24, Sherman2023, Drescher2023, janssenConfinement2020]. In this condensed matter context, the gapless and critical ground state is referred to as the $U(1)$ DSL.
@@
From the field-theoretical viewpoint, the basic structure of QED3 CFT is understood—one knows the set of primary operators and rough estimates of their scaling dimensions [di2017scaling, chester2016towards, Albayrak2022]. Such quantitative information about the scaling dimensions of fermion bilinear operators is important, since the scaling dimensions govern the asymptotic decay of correlation functions and diagnose the conformal nature of the DSL. In particular, the spin, bond, and dimer correlations correspond to different fermion bilinears that transform in the adjoint representation of the emergent $SU(4)$ symmetry for $N_{f}=2$, and are expected to share the same scaling exponent [rantner2002spin, Hermele2005]. Establishing this universality and comparing with analytical approaches such as large-$N_{f}$ expansion or $\epsilon$-expansion [di2017scaling, chester2016towards] is crucial to confirm the conformal fixed point.
@@
On the condensed matter experimental side, signatures of proximate DSLs have been reported in frustrated magnets, including triangular lattice antiferromagnets YbZn2GaO5 [BagEvidence24, WuSpinDynamics25] and A-YbSe2 delafossites [ScheieKYbSe2, ScheieNaYbSe2], as well as kagomé antiferromagnets [zengPossible2022, zengSpectral2024, hanSpin2024, suetsuguEmergent2024, suetsuguGapless2024, jeonOne2024, zhengUnconventional2025]. For example, the recently synthesized kagomé compound YCu3(OD)6Br2[Br1-x(OD)x] shows a clear $T^{2}$ contribution in specific heat [zengPossible2022], inelastic neutron scattering reveals a continuum consistent with Dirac cones [zengSpectral2024, hanSpin2024], and magnetic torque detects unconventional oscillations near the 1/9 plateau [zhengUnconventional2025]. These results point toward DSL behavior, but a quantitative comparison requires reliable knowledge of the scaling of correlation functions.
@@
On the computational side, progress has been hindered by the absence of efficient and unbiased methods. Determinant quantum Monte Carlo (DQMC) has been applied to finite-size lattice QED3 [XYXuU12018, wangDynamics2019, janssenConfinement2020], but its computational cost scales as $O(N_{\tau}V_{s}^{3})$, with $N_{\tau}$ the imaginary-time extent and $V_{s}$ the spatial volume. This cubic scaling, together with slow dynamics of local gauge updates, restricts simulations to small lattices and low precision. In critical phases such as the DSL, DQMC also suffers from severe autocorrelation effects [panSelf2025], further reducing efficiency. As a result, DQMC has not been able to access sufficiently large scales to settle questions about operator scaling dimensions or to confirm the conformal properties of QED3.
@@
Hybrid quantum Monte Carlo (HQMC) offers an appealing alternative [Scalettar1987, Duane1987, Gupta1988, Beyl2018]. Unlike local-update DQMC, HQMC updates gauge field configurations globally using Hamiltonian dynamics, while the pseudofermion representation replaces explicit determinant evaluations by iterative solutions of linear systems. In principle this reduces the complexity from $O(N_{\tau}V_{s}^{3})$ to $O(N_{\tau}V_{s})$, and is inherently well suited to parallelization. Yet in earlier condensed matter applications, HQMC faced difficulties such as poor matrix conditioning and diverging iteration counts [Scalettar1987, White1988], which limited its practical advantages. More recently, HQMC has been successfully applied to spin-fermion models [patel2024strange, lunts2023non], where optimizations including preconditioned conjugate gradient solvers, matrix-free multiplications, and GPU acceleration enabled large-scale simulations. With these computational improvements, the non-Fermi-liquid scaling in the fermionic self-energy and bosonic propagators (space-time correlation functions) are observed. These encouraging developments motivate a dedicated HQMC approach to lattice QED3.
@@
Inspired by this progress, in this paper we apply HQMC to simulate a $U(1)$ gauge field model coupled to $N_{f}=2$ fermions in (2+1)d, fully accelerated on GPUs. We scale the simulation size up to $N_{\tau}\times L\times L=660\times 66\times 66$, which represents a major improvement over previous DQMC studies of this model [XYXuU12018, wangDynamics2019, janssenConfinement2020] (typically limited to $200\times 20\times 20$), and exceeds by more than a factor of three the largest HQMC applications in spin-fermion models [patel2024strange, lunts2023non]. This advance is achieved through three problem-specific GPU optimizations:
@@
With these optimizations, the computational complexity is reduced from DQMC’s $O(N_{\tau}V_{s}^{3})$ scaling to nearly linear $O(N_{\tau}V_{s})$, allowing us to reach an unprecedented regime of system sizes. At this scale, we are able to investigate the correlation functions of various fermion bilinear operators and conserved currents with asymptotic accuracy. We find that spin and bond correlations share the same scaling dimension in the thermodynamic limit, consistent with their identification as adjoint bilinears of the emergent $SU(4)$ symmetry. The extracted scaling dimension $\Delta_{\rm adj}$ lies in the range 1.751.751.951.95, in agreement with the most recent analytical $\epsilon$-expansion results [di2017scaling]. We further observe that flux–flux correlations are governed by the conserved current dimension $\Delta_{J}=2$, leading to asymptotic long-time decay $\sim 1/\tau^{4}$. These results provide new, unbiased evidence for the conformal nature of the $U(1)$ DSL in QED3. Beyond these physical findings, our computational developments establish a scalable framework to explore DSL stability and its transitions to symmetry-breaking phases with significantly reduced computational cost.
@@ I Results
The models we study are quantum electrodynamics (QED3) defined on a cubic space-time lattice (square spatial lattice), coupled with two-flavor fermions, as shown in Fig. 1 [XYXuU12018, wangDynamics2019]. Here, we start with the action of these lattice QED3 models: $S=S_{F}+S_{B}=\int_{0}^{\beta}d\tau\left(L_{F}+L_{B}\right)$, where $\beta$ is the inverse temperature and $\tau$ is the imaginary time. The first term
@@
In this paper, we focus on a dubbed non-compact non-compact $U(1)$ gauge field described by the Lagrangian
$L_{B}^{\textmd{nc}}=\frac{1}{J\Delta\tau^{2}}\sum_{\langle ij\rangle}\left[\phi_{ij,\tau+1}-\phi_{ij,\tau}\right]^{2}+K\sum_{\square}\cos(\operatorname{curl}\phi),$ (2)
where $J$ and $K$ are the coupling constants, $\Delta\tau=0.1$ is the imaginary time step introduced in the Trotter decomposition [XYXuU12018], and $\textmd{curl}\phi=\sum_{ij\in\textmd{edge}(\Box)}\phi_{ij}$ is a discretized curvature of the gauge field on a spacial plaquette as shown in Fig. 1. The name non-compact comes from the fact that the gauge flux, defined as the discretized curvature of the gauge field [Fiebig1990, degrand1980topological] on a face of the space-time cube illustrated in Fig. 1, does not have $2\pi$-periodicity, as opposed to the compact Lagrangian, e.g. $L_{B}^{\textmd{c}}=\frac{1}{J\Delta\tau^{2}}\sum_{\langle ij\rangle}\left[1-\cos\left(\phi_{ij,\tau+1}-\phi_{ij,\tau}\right)\right]+K\sum_{\square}\cos(\operatorname{curl}\phi)$, where the gauge flux does have $2\pi$-periodicity (see methods section). As a result, the non-compact model does not contain gauge monopoles, which gives a clean background to study the deconfined U(1) Dirac spin liquid phase in this model [XYXuU12018].
@@
Figure 2: The speedup achieved by customized CUDA kernel and CUDA Graph, and the computational complexity of HQMC. (a) shows the speedup achieved by customized CUDA kernel for different system sizes $L$, with $V_{s}=L^{2}$ and $N_{\tau}=10L$. The blue axis on the left represents the latency of the Monte Carlo sweep which is average time cost per generated sample. The light blue line represents the baseline performance, where the PyTorch sparse-matrix-vector multiplication is applied to compute $M^{\dagger}Mv$ and $\tilde{O}^{-1}v$. The dark blue line is the latency of the customized CUDA kernel, where the computations $M^{\dagger}Mv$ and $\tilde{O}^{-1}v$ are replaced with optimized CUDA kernel. The red dashed line is the latency ratio, which shows the speedup between the two implementations for each system size. (b) shows the speedup achieved by CUDA Graph, for different system sizes $L$. The light blue line represents the baseline performance with CUDA Graph turned off, while the dark blue line is the latency when CUDA Graph is turned on. The red dashed line is the latency ratio, which shows the speedup between the two implementations for each system size. (c) shows the scaling behavior of the HQMC sample generation latency and HQMC measurement latency using stochastic estimation (SE), as functions of $N_{\tau}V_{s}\propto L^{3}$, where $N_{\tau}=\beta/\Delta_{\tau}=10L$. The latency (second per sample) is defined as the time spent to generate a gauge sample, or the time to compute a fermionic observable from a gauge sample for SE (the SE latency here is tested on spin-spin correlation using $N_{\textmd{rv}}=40$). The blue and orange line are the linear fitting of five large-size points. They show that the HQMC algorithm has a time complexity of $O((L^{3})^{0.9})$, while the stochastic estimation has a time complexity of $O((L^{3})^{1.3})$. As a contrast, the DQMC has a time complexity of $O(N_{\tau}V_{s}^{3})=O(L^{7})$, which is plotted as the green dashed line. The starting point of the green dashed line is $0.38$ second per sweep at $L^{3}=10^{3}$, obtained from DQMC benchmark.
@@
First, our customized CUDA kernel is designed to effectively exploit the sparsity pattern of the large matrices, and puts the reusable matrix entries in the shared memory of GPU, which is also called local caching [lunts2023non]. In this way, the communication overhead within GPU is significantly reduced, and a 2.9 - 3.8 times speedup is achieved compared to naive GPU implementation, which is shown in Fig.2(a). This local caching is an advanced optimization technique in large language model system [dao2022flashattention], and now finds its application in broader scientific computing community [lunts2023non, patel2024strange].
Second, CUDA graph is another important optimization technique in our simulation. This technique was not introduced until very recently, has been applied in a variety of large language model systems [kwon2023efficient, zheng2024sglang], and just starts to find its application in the broader scientific computing field. This technique compiles a series of CUDA kernel launches into a single CUDA graph, which only needs to be launched once. Thus, the large amount of repetitive CUDA kernel launching overhead is spared (see methods section for details). As illustrated in Fig.2(b), this CUDA graph optimization achieves a speed up of around 2-5 times on moderate lattice sizes (around $L<30$) and 1-2 times on large system sizes, compared to the baseline where CUDA Graph is turned off. The technical details are presented in Methods section.
Third, finding an efficient preconditioner is challenging. There is no universal way to compute it, and is still an ongoing research [yamazaki2009high, beyl2018revisiting] on its own. In addition, when scaling up the simulation to $N_{\tau}\times L^{2}=660\times 66^{2}$, we have to face extra challenge of memory bottleneck when computing the preconditioner, which could push the peak memory up to thousands of gigabytes. However, specific to the problem studied here, we find a proper preconditioner that successfully addresses all the above challenges. The key relies on the two-fold generalizability in our preconditioner. The first is that the preconditioner obtained by one gauge configuration can be globally applied on other generic gauge configurations, without the need of local recomputation. The second is that the entries of the preconditioner at large lattice sizes can be directly inferred from the small lattice sizes, without any computation (see Methods section). These findings can be applied to the study of other problems involving the gauge-field-connected fermion hopping Lagrangian like Eq. (1).
@@
Here, the spin operator is defined as $S^{\alpha\beta}(i)=\psi_{i\alpha}^{\dagger}\psi_{i\beta}-\frac{1}{N_{f}}\delta_{\alpha\beta}\sum_{\gamma}\psi_{i\gamma}^{\dagger}\psi_{i\gamma}$, and the bond operator along the nearest-neighbor bond in $\hat{x}$ direction is defined as $B_{i}=\sum_{\alpha}(\psi^{\dagger}_{i,\alpha}e^{i\phi_{i,i+\hat{x}}}\psi_{i+\hat{x},\alpha}+\text{h.c.})$. $\alpha,\beta\in\left\{\uparrow,\downarrow\right\}$ are spin flavor indices. We have written the spin operators as antisymmetric self-conjugate representation [Hermele2004]. They are related to the normal spin operators definition as $S^{\uparrow\downarrow}=S^{+},S^{\downarrow\uparrow}=S^{-},S^{\uparrow\uparrow}=S^{z},S^{\downarrow\downarrow}=-S^{z}$.
Other than the gauge invariant fermionic observables, we also measure the correlation of the magnetic flux operator $\theta_{\square}:=\textmd{curl}\phi=\sum_{ij\in\textmd{edge}(\Box)}\phi_{ij}$, where $\phi_{ij}=-\phi{ji}$ and the summation is along the positive directions shown in Fig. 1. We focus on the imaginary time dynamics. Since, in the static limit ($J=0$) of the QED3, the ground energy magnetic configuration is known to be $\pi$-flux on a square lattice [lieb1994flux]. When $J$ is turned on, the fluctuation of magnetic flux is centered also around $\pi$-flux. So, to study the correlation of the magnetic fluxes, we measure the correlation of its sine value instead, in order to have a clean background, i.e. $\langle\sin\theta_{\square}\rangle=0$. We define
@@
Next, the CFT operator scaling dimension are readily extracted from the correlation functions. As introduced before, the DSL with non-compact QED3 has been analyzed extensively and is believed to describe a scale-invariant CFT. Consequently, power law behavior is expected for all gauge-invariant observables. In our setting of $N_{f}=2$, the CFT is expected to have emergent $SU(4)$ symmetry (with two cones in the BZ), one expects the scaling dimension of the fermion bilinears or mass terms to be the following: for a singlet $\Delta_{s}\approx 2.4$ and for a set of adjoints $\Delta_{\text{adj}}\in(1.75,1.95)$, where there are 15 of them ranging over the $SU(4)$ generators. These scaling dimensions are analytically obtained by the $\epsilon$-expansion from $d=4-2\epsilon$ to $d=3$, in one of the most up-to-date computations[di2017scaling]. We note that, the spin operator and the bond operator mentioned above in Eqs. (4) and (5) are among the $15$ adjoint fermion bilinear terms and consequently they are expected to share the same power-law ($\sim 1/r^{2\Delta_{\text{adj}}}$) when we compute their correlation functions [Hermele2005, XYXuU12018] and it is the indeed the case as we show in the next section.
@@
Moreover, in the $U(1)$ DSL phase, one expects that the flux-flux correlations is dictated by the scaling dimension of the conserved current $\Delta_{J}=2$, which leads to the long-time correlations of $C_{\textmd{flux}}(\tau)\sim\tau^{-4}$. In contrast, for a non-compact $U(1)$ pure free-photon theory without fermion coupling, which is non-conformal, the flux correlation decays with a different power, $C_{\textmd{flux}}(\tau)\sim\tau^{-3}$ [chenEmergent2025]. From our HQMC simulation, the $\tau^{-4}$ behavior of flux correlation is clearly seen at large space-time volume.

### v4

@@ Introduction
The emergence of Dirac spin liquid (DSL), i.e. an emergent $U(1)$ gauge field coupled to gapless relativistic spinons in (2+1)d [31, 6, 32, 50, 51, 67, 63, 43, 65, 57, 37], or, in high-energy physics terms, a deconfined phase in 3D quantum electrodynamics (QED3) with $N_{f}$ flavors of massless fermions [23, 30, 24, 5, 13, 17, 39, 3], has attracted continuous attention from condensed matter and high-energy physicists over the past decades. For sufficiently large $N_{f}$, QED3 is expected to flow to a conformal field theory (CFT) in the infrared [39, 13], containing a variety of scaling operators including fermion bilinear currents, mass operators, and monopole operators. Such a critical state has been argued to emerge in frustrated spin systems, where no magnetic order appears and the spinons acquire Dirac dispersions, and their interactions are mediated by fluctuating $U(1)$ gauge fields [28, 31, 6, 32, 50, 51, 35, 1, 73, 34, 36, 29, 25, 53, 67, 63, 59, 33, 22, 44, 43, 65, 57, 58, 18, 37]. In this condensed matter context, the gapless and critical ground state is referred to as the $U(1)$ DSL.
@@
From the field-theoretical viewpoint, the basic structure of QED3 CFT is understood—one knows the set of primary operators and rough estimates of their scaling dimensions [17, 13, 3]. Such quantitative information about the scaling dimensions of fermion bilinear operators is important, since the scaling dimensions govern the asymptotic decay of correlation functions and diagnose the conformal nature of the DSL. In particular, the spin, bond, and dimer correlations correspond to different fermion bilinears that transform in the adjoint representation of the emergent $SU(4)$ symmetry for $N_{f}=2$, and are expected to share the same scaling exponent [52, 32]. Establishing this universality and comparing with analytical approaches such as large-$N_{f}$ expansion or $\epsilon$-expansion [17, 13] is crucial to confirm the conformal fixed point.
@@
On the condensed matter experimental side, signatures of proximate DSLs have been reported in frustrated magnets, including triangular lattice antiferromagnets YbZn2GaO5 [7, 66] and A-YbSe2 delafossites [55, 56], as well as kagomé antiferromagnets [69, 70, 27, 61, 60, 38, 71]. For example, the recently synthesized kagomé compound YCu3(OD)6Br2[Br1-x(OD)x] shows a clear $T^{2}$ contribution in specific heat [69], inelastic neutron scattering reveals a continuum consistent with Dirac cones [70, 27], and magnetic torque detects unconventional oscillations near the 1/9 plateau [71]. These results point toward DSL behavior, but a quantitative comparison requires reliable knowledge of the scaling of correlation functions.
@@
On the computational side, progress has been hindered by the absence of efficient and unbiased methods. Determinant quantum Monte Carlo (DQMC) has been applied to finite-size lattice QED3 [67, 63, 37], but its computational cost scales as $O(N_{\tau}V_{s}^{3})$, with $N_{\tau}$ the imaginary-time extent and $V_{s}$ the spatial volume. This cubic scaling, together with slow dynamics of local gauge updates, restricts simulations to small lattices and low precision. In critical phases such as the DSL, DQMC also suffers from severe autocorrelation effects [46], further reducing efficiency. As a result, DQMC has not been able to access sufficiently large scales to settle questions about operator scaling dimensions or to confirm the conformal properties of QED3.
@@
Hybrid quantum Monte Carlo (HQMC) offers an appealing alternative [54, 19, 26, 8]. Unlike local-update DQMC, HQMC updates gauge field configurations globally using Hamiltonian dynamics, while the pseudofermion representation replaces explicit determinant evaluations by iterative solutions of linear systems. In principle this reduces the complexity from $O(N_{\tau}V_{s}^{3})$ to $O(N_{\tau}V_{s})$, and is inherently well suited to parallelization. Yet in earlier condensed matter applications, HQMC faced difficulties such as poor matrix conditioning and diverging iteration counts [54, 64], which limited its practical advantages. More recently, HQMC has been successfully applied to spin-fermion models [47, 42], where optimizations including preconditioned conjugate gradient solvers, matrix-free multiplications, and GPU acceleration enabled large-scale simulations. With these computational improvements, the non-Fermi-liquid scaling in the fermionic self-energy and bosonic propagators (space-time correlation functions) are observed. These encouraging developments motivate a dedicated HQMC approach to lattice QED3.
@@
Inspired by this progress, in this paper we apply HQMC to simulate a $U(1)$ gauge field model coupled to $N_{f}=2$ fermions in (2+1)d, fully accelerated on GPUs. We scale the simulation size up to $N_{\tau}\times L\times L=660\times 66\times 66$, which represents a major improvement over previous DQMC studies of this model [67, 63, 37] (typically limited to $200\times 20\times 20$), and exceeds by more than a factor of three the largest HQMC applications in spin-fermion models [47, 42]. This advance is achieved through three problem-specific GPU optimizations:
@@
With these optimizations, the computational complexity is reduced from DQMC’s $O(N_{\tau}V_{s}^{3})$ scaling to nearly linear $O(N_{\tau}V_{s})$, allowing us to reach an unprecedented regime of system sizes. At this scale, we are able to investigate the correlation functions of various fermion bilinear operators and conserved currents with asymptotic accuracy. We find that spin and bond correlations share the same scaling dimension in the thermodynamic limit, consistent with their identification as adjoint bilinears of the emergent $SU(4)$ symmetry. The extracted scaling dimension $\Delta_{\rm adj}$ lies in the range 1.751.751.951.95, in agreement with the most recent analytical $\epsilon$-expansion results [17]. We further observe that flux–flux correlations are governed by the conserved current dimension $\Delta_{J}=2$, leading to asymptotic long-time decay $\sim 1/\tau^{4}$. These results provide new, unbiased evidence for the conformal nature of the $U(1)$ DSL in QED3. Beyond these physical findings, our computational developments establish a scalable framework to explore DSL stability and its transitions to symmetry-breaking phases with significantly reduced computational cost.
@@ I Results
The models we study are quantum electrodynamics (QED3) defined on a cubic space-time lattice (square spatial lattice), coupled with two-flavor fermions, as shown in Fig. 1 [67, 63]. Here, we start with the action of these lattice QED3 models: $S=S_{F}+S_{B}=\int_{0}^{\beta}d\tau\left(L_{F}+L_{B}\right)$, where $\beta$ is the inverse temperature and $\tau$ is the imaginary time. The first term
@@
In this paper, we focus on a dubbed non-compact $U(1)$ gauge field described by the Lagrangian
$L_{B}^{\textmd{nc}}=\frac{1}{J\Delta\tau^{2}}\sum_{\langle ij\rangle}\left(\phi_{ij,\tau+1}-\phi_{ij,\tau}\right)^{2}+K\sum_{\square}\cos(\operatorname{curl}\phi),$ (2)
where $J$ and $K$ are the coupling constants, $\Delta\tau=0.1$ is the imaginary time step introduced in the Trotter decomposition [67], and $\textmd{curl}\phi=\sum_{ij\in\textmd{edge}(\Box)}\phi_{ij}$ is a discretized curvature of the gauge field on a spacial plaquette as shown in Fig. 1. The name non-compact comes from the fact that the gauge flux, defined as the discretized curvature of the gauge field [23, 16] on a face of the space-time cube illustrated in Fig. 1, does not have $2\pi$-periodicity, as opposed to the compact Lagrangian, e.g. $L_{B}^{\textmd{c}}=\frac{1}{J\Delta\tau^{2}}\sum_{\langle ij\rangle}\left[1-\cos\left(\phi_{ij,\tau+1}-\phi_{ij,\tau}\right)\right]+K\sum_{\square}\cos(\operatorname{curl}\phi)$, where the gauge flux does have $2\pi$-periodicity (see methods section). As a result, the non-compact model does not contain gauge monopoles, which gives a clean background to study the deconfined U(1) Dirac spin liquid phase in this model [67].
@@
Figure 2: The speedup achieved by customized CUDA kernel and CUDA Graph, and the computational complexity of HQMC. Panel (a) shows the speedup achieved by customized CUDA kernel for different system sizes $L$, with $V_{s}=L^{2}$ and $N_{\tau}=10L$. The blue axis on the left represents the latency of the Monte Carlo sweep which is average time cost per generated sample. The light blue line represents the baseline performance, where the PyTorch sparse-matrix-vector multiplication is applied to compute $M^{\dagger}Mv$ and $\tilde{O}^{-1}v$. The dark blue line is the latency of the customized CUDA kernel, where the computations $M^{\dagger}Mv$ and $\tilde{O}^{-1}v$ are replaced with optimized CUDA kernel. The red dashed line is the latency ratio, which shows the speedup between the two implementations for each system size. (b) shows the speedup achieved by CUDA Graph, for different system sizes $L$. The light blue line represents the baseline performance with CUDA Graph turned off, while the dark blue line is the latency when CUDA Graph is turned on. The red dashed line is the latency ratio, which shows the speedup between the two implementations for each system size. (c) shows the scaling behavior of the HQMC sample generation latency and HQMC measurement latency using stochastic estimation (SE), as functions of $N_{\tau}V_{s}\propto L^{3}$, where $N_{\tau}=\beta/\Delta_{\tau}=10L$. The latency (second per sample) is defined as the time spent to generate a gauge sample, or the time to compute a fermionic observable from a gauge sample for SE (the SE latency here is tested on spin-spin correlation using $N_{\textmd{rv}}=40$). The blue and orange line are the linear fitting of five large-size points. They show that the HQMC algorithm has a time complexity of $O((L^{3})^{0.9})$, while the stochastic estimation has a time complexity of $O((L^{3})^{1.3})$. As a contrast, the DQMC has a time complexity of $O(N_{\tau}V_{s}^{3})=O(L^{7})$, which is plotted as the green dashed line. The starting point of the green dashed line is $0.38$ second per sweep at $L^{3}=10^{3}$, obtained from DQMC benchmark.
@@
First, our customized CUDA kernel is designed to effectively exploit the sparsity pattern of the large matrices, and puts the reusable matrix entries in the shared memory of GPU, which is also called local caching [42]. In this way, the communication overhead within GPU is significantly reduced, and a 2.9 - 3.8 times speedup is achieved compared to naive GPU implementation, which is shown in Fig.2(a). This local caching is an advanced optimization technique in large language model system [15], and now finds its application in broader scientific computing community [42, 47].
Second, CUDA graph is another important optimization technique in our simulation. This technique was not introduced until very recently, has been applied in a variety of large language model systems [40, 72], and just starts to find its application in the broader scientific computing field. This technique compiles a series of CUDA kernel launches into a single CUDA graph, which only needs to be launched once. Thus, the large amount of repetitive CUDA kernel launching overhead is spared (see methods section for details). As illustrated in Fig.2(b), this CUDA graph optimization achieves a speed up of around 2-5 times on moderate lattice sizes (around $L<30$) and 1-2 times on large system sizes, compared to the baseline where CUDA Graph is turned off. The technical details are presented in Methods section.
Third, finding an efficient preconditioner is challenging. There is no universal way to compute it, and is still an ongoing research [68, 9] on its own. In addition, when scaling up the simulation to $N_{\tau}\times L^{2}=660\times 66^{2}$, we have to face extra challenge of memory bottleneck when computing the preconditioner, which could push the peak memory up to thousands of gigabytes. However, specific to the problem studied here, we find a proper preconditioner that successfully addresses all the above challenges. The key relies on the two-fold generalizability in our preconditioner. The first is that the preconditioner obtained by one gauge configuration can be globally applied on other generic gauge configurations, without the need of local recomputation. The second is that the entries of the preconditioner at large lattice sizes can be directly inferred from the small lattice sizes, without any computation (see Methods section). These findings can be applied to the study of other problems involving the gauge-field-connected fermion hopping Lagrangian like Eq. (1).
@@
Here, the spin operator is defined as $S^{\alpha\beta}(i)=\psi_{i\alpha}^{\dagger}\psi_{i\beta}-\frac{1}{N_{f}}\delta_{\alpha\beta}\sum_{\gamma}\psi_{i\gamma}^{\dagger}\psi_{i\gamma}$, and the bond operator along the nearest-neighbor bond in $\hat{x}$ direction is defined as $B_{i}=\sum_{\alpha}(\psi^{\dagger}_{i,\alpha}e^{i\phi_{i,i+\hat{x}}}\psi_{i+\hat{x},\alpha}+\text{h.c.})$. $\alpha,\beta\in\left\{\uparrow,\downarrow\right\}$ are spin flavor indices. We have written the spin operators as antisymmetric self-conjugate representation [31]. They are related to the normal spin operators definition as $S^{\uparrow\downarrow}=S^{+},S^{\downarrow\uparrow}=S^{-},S^{\uparrow\uparrow}=S^{z},S^{\downarrow\downarrow}=-S^{z}$.
Other than the gauge invariant fermionic observables, we also measure the correlation of the magnetic flux operator $\theta_{\square}:=\textmd{curl}\phi=\sum_{ij\in\textmd{edge}(\Box)}\phi_{ij}$, where $\phi_{ij}=-\phi{ji}$ and the summation is along the positive directions shown in Fig. 1. We focus on the imaginary time dynamics. Since, in the static limit ($J=0$) of the QED3, the ground energy magnetic configuration is known to be $\pi$-flux on a square lattice [41]. When $J$ is turned on, the fluctuation of magnetic flux is centered also around $\pi$-flux. So, to study the correlation of the magnetic fluxes, we measure the correlation of its sine value instead, in order to have a clean background, i.e. $\langle\sin\theta_{\square}\rangle=0$. We define
@@
Next, the CFT operator scaling dimension are readily extracted from the correlation functions. As introduced before, the DSL with non-compact QED3 has been analyzed extensively and is believed to describe a scale-invariant CFT. Consequently, power law behavior is expected for all gauge-invariant observables. In our setting of $N_{f}=2$, the CFT is expected to have the emergent $SU(4)$ symmetry [17]. Here, $N_{f}=2$ counts the number of four-component Dirac fermion flavor. More specifically, the spin $\{\uparrow,\downarrow\}$ degrees of freedom go into the four-component Dirac fermion, while the $N_{f}=2$ correspond the two spin-degenerate Dirac cones in the first Brillouin Zone. One expects the scaling dimension of the fermion bilinears or mass terms to be the following: for a singlet $\Delta_{s}\approx 2.4$ and for a set of adjoints $\Delta_{\text{adj}}\in(1.75,1.95)$, where there are 15 of them ranging over the $SU(4)$ generators; these scaling dimensions are analytically obtained by the $\epsilon$-expansion from $d=4-2\epsilon$ to $d=3$, in one of the most up-to-date computations [17]. We note that, the spin operator and the bond operator mentioned above in Eqs. (4) and (5) are among the $15$ adjoint fermion bilinear terms and consequently they are expected to share the same power-law ($\sim 1/r^{2\Delta_{\text{adj}}}$) when we compute their correlation functions [32, 67] and it is the indeed the case as we show in the next section.
@@
Moreover, in the $U(1)$ DSL phase, one expects that the flux-flux correlations is dictated by the scaling dimension of the conserved current $\Delta_{J}=2$, which leads to the long-time correlations of $C_{\textmd{flux}}(\tau)\sim\tau^{-4}$. In contrast, for a non-compact $U(1)$ pure free-photon theory without fermion coupling, which is non-conformal, the flux correlation decays with a different power, $C_{\textmd{flux}}(\tau)\sim\tau^{-3}$ [11]. From our HQMC simulation, the $\tau^{-4}$ behavior of flux correlation is seen at large space-time volume.